Circular No. 9090

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 U6 (LINEAR) An apparently asteroidal object discovered by the LINEAR survey

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to show cometary appearance by other CCD astrometrists. T. Kryachko, Moscow, writes that images taken by B. Satovski and himself with a 0.3-m f/7.7 reflector at the Zelenchukskaya Station of Engelhardt Observatory on Oct. 28.1 UT show a diffuse coma of diameter 8′′ and a 30′′ tail in p.a. 290^{o}. F. Fratev (Zvezdno Obshtestvo Observatory, Plana, Bulgaria, 0.35-m f/3.4 reflector, Oct. 29.1) reports that fifteen stacked 40-s unfiltered exposures show a diffuse coma nearly 10′′ in diameter and a tail ∼ 30′′ long in p.a. 305^{o}. J. L. Montani notes that ≈ 120-s drift-scan images taken by T. H. Bressi on Oct. 29.5 with the Spacewatch 1.8-m reflector at Kitt Peak show a straight, fan-shaped, 80′′ tail in p.a. 288^{o}; the tail is 15′′ wide at 80′′ from the nucleus.

2009 UT  α_{2000}  δ_{2000}  Mag.
Oct. 27.43474  9^{h}44^{m}36.^{s}35  +13^{o}04′54′′.7  18.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-U126.

T = 2009 Aug. 26.433 TT  ω = 319.^{o}286
Ω = 115.014 }2000.0
q = 1.61891 AU  i = 14.334

COMETS C/2009 J11, C/2009 J12, AND C/2009 K6 (SOHO) Further to IAUC 9088, two additional near-sun presumed comets have

been found on SOHO website images. C/2009 J11, a Kreutz sungrazer, was tiny and stellar in appearance in C3 images (peak mag ∼ 7–7.5), but slightly diffuse with a hint of a very faint tail in C2 images. C/2009 J12 (a Marsden-group member; mag ∼ 7.5–8) and C/2009 K6 (also a Kreutz sungrazer; mag ∼ 7) were also stellar in appearance.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 J11 May 13.221 3^{h}23.^{m}9 +16^{o}24′ C3/2 BZ 2009-N49
C/2009 J12  15.579 3 34.5 +19 23 C2  RK 2009-N49
C/2009 K6  17.229 3 38.0 +17 48 C2  BZ 2009-N49
2009 October 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green